# Imperceptible Protection against Style Imitation from Diffusion Models

### v2

@@ Appendix C Experimental Settings
User study. We conducted a user study in the form of an A/B test with a reference image as a benchmark. A total of 44 participants were involved, tasked with determining the better-quality protected image (given the reference original artwork) and identifying the lower-quality generated image (given a reference DreamBooth (Ruiz et al. 2023) generated image). Each participant was asked to vote on 8 questions for PhotoGuard (Salman et al. 2023) and another 9 for AdvDM (Liang et al. 2023).
Appendix D Discussions and Analyses
Full vs. partial protection. Extending Figure 2, in Figure 11, we compare the original artwork, and the personalization results of unprotected, partially protected, and fully protected artworks. Partial protection simulates hard restrictions, whereas full protection mimics soft restrictions. To create partially protected images, we apply protection to the center (Figure 11, top) or the facial region (Figure 11, bottom), as indicated by the red boxes. Both the unprotected and partially protected images cannot prevent against style imitation from the personalization method. In contrast, full protection (soft restriction) successfully prevents this.
We also conduct a quantitative comparison between partial and full protection with FID on the painting dataset. In this experiment, the partial protection achieves 260.2, while full protection scores 299.1, demonstrating the higher efficacy of full protection. This empirical evidence led us to propose soft restriction-based JND perceptual protection.
JNDs. In this section, we further analyze the protection performance of JND-based perception map depicted in Figure 3 (right). When compared to the baseline, perceptual map improves image quality (DISTS) across the board, but it also leads to a compromise in protection performance (FID). Among the JNDs, LA demonstrates the best protection performance. We conjecture that in many artworks, the majority of areas fall high or low-luminance, thereby maintaining perturbations strength high across extensive regions. However, as shown in upper image, even simple textures (e.g. sky) can have strong perturbations, placing LA at the lower image quality. On the other hand, CSF, Std, and Entropy generally apply high perturbations only to specific areas, such as edges, resulting in most regions being not fully protected and consequently, causing a huge degradation in protection performance. CM’s protection intensity is also determined by spatial changes but this covers more detailed local regions (see fields in upper image) and also being based on contrast, leading to both high quality and effective protection.
Ablation study. In Table 2, one might expect that the JND (PAP with JND) maintains protection scores while improving image quality. Indeed, the random mask (PAP without JND) constrains the protection intensity in a similar level to the JND-based map. This is because both maps are normalized between 0 and 1. As a result, they yield similar perturbation magnitudes, leading to a comparable image quality. Nonetheless, the JND-based perceptual map prioritizes less sensitive regions for perturbation while reducing it in highly sensitive areas. Despite the similar protection magnitudes, the JND-based PAP achieves more effective protection. This hints at the possibility that the JND maps are helpful in finding the areas that are important for style protection. For instance, in areas with complex textures, the PAP applies more perturbations than its non-JND counterpart. However, these perturbations are typically imperceptible to humans, thereby ensuring perceptually acceptable image quality.
In Table 4, we show additional analysis by omitting certain components from the full Impasto framework. Specifically, we assess the impact of 1) removing the JND-based perception map, which results in initializing $\mathbf{M}$ with random masks (w/o JND), and 2) excluding the proposed perception-aware protection (PAP), thereby relying solely on the perceptual constraint bank (w/o PAP). For the w/o JND case, in contrast to Table 2, we start with a full-component model and remove only the JND part. The results indicate that discarding the JNDs leads to a decrease in protection performance while maintaining similar image quality, corroborating the findings in Table 2. When PAP is not applied, there is a slight increase in protection performance (FID: 279.3 $\rightarrow$ 281.1) with significant compromise on image quality. These observations suggest that the integration of PAP alongside a JND-based perceptual map is crucial for achieving an imperceptible style protection framework that effectively balances protection performance with image quality.
Table 8: Quantitative comparison of protection methods w/ and w/o Impasto, both selected for their comparable protection performance. Impasto markedly elevates the protected images’ quality while maintaining protection efficacy.
@@ Appendix D Discussions and Analyses
DISTS ($\downarrow$) LPIPS ($\downarrow$) TOPIQ ($\uparrow$) NIQE ($\uparrow$) BRISQUE ($\uparrow$) FID ($\uparrow$)
Painting Anti-DB 0.151 (+0.000) 0.081 (+0.000) 0.876 (+0.000) 3.865 13.63 266.1
Mist 0.167 (+0.000) 0.100 (+0.000) 0.846(+0.000) 4.052 13.96 272.2
Cartoon Anti-DB 0.260 (+0.000) 0.154 (+0.000) 0.700(+0.000) 4.437 15.48 160.6
Mist 0.256 (+0.000) 0.160 (+0.000) 0.709(+0.000) 4.597 10.86 158.7
Varying refinement interval step. In this analysis, we explore the effect of varying the refinement interval step $P$ on instance-wise refinement (IWR) and difficulty-aware protection (DAP) to evaluate changes in protection performance (Table 5). We observe that even as $P$ increases from 1 to 4 (i.e., with less frequent refinements), both image quality and protection capability remain relatively stable. This stability can be attributed to the initial use of a JND-based perceptual map $\mathcal{M}$, which requires only minor adjustments during the IWR step. However, when $P$ is increased to 10, a significant reduction in protection performance is observed, indicating that such a large interval is insufficient for optimal refinement. Therefore, we set $P=4$ as the default, balancing performance efficacy with the number of optimization steps.
Table 9: Comparison on the blur countermeasure of PhotoGuard with and without Impasto at a similar level of protected image quality.
FID ($\uparrow$) PhotoGuard + Impasto
Clean 274.4 279.3
Blur 273.5 274.6
Countermeasures. As discussed earlier, Impasto achieves similar protection performance against JPEG compression and Gaussian noise countermeasures but shows slight degradation under Gaussian blur. The reason for this phenomenon is that the blurring effect removes most high-frequencies, making weak perturbations less effective. We note that if baseline protection models are applied with weak protection strength (to match the protected images’ perceptual quality to ours), they also exhibit degraded protection performance.
To demonstrate this, we match the protected image quality of PhotoGuard to ours and compare the protection performance in both clean (no countermeasure) and blur scenarios (Table 9). In this case, which with similar imperceptibility, Impasto shows better performance under blur, implying that such degradation is an inevitable limitation of using weaker perturbations rather than a flaw in our model design. Note that the performance gap narrows under blur because Impasto applies very weak perturbations to more noticeable regions, potentially causing minor vulnerabilities. Nevertheless, in all scenarios—whether clean or under countermeasures—Impasto demonstrates an improved trade-off between protected image quality and protection performance, highlighting its practical effectiveness.
@@
Protection performance comparison. Here, unlike Table 1, which compares the protected images’ quality under similar protection performance, we compare the protection performance while fixing the image quality. In Table 6, Impasto improves protection performance compared to the baseline in most metrics and scenarios. This was also shown in the quality-protection trade-off (Figure 7). These results indicate that attaching Impasto to the protection baselines enables them to be better protection methods.
Impasto in other domain. Although Impasto is initially proposed to prevent style imitation, its applicability can be extended beyond other domains or applications. To validate this, we conduct protection on two facial datasets, CelebA-HQ (Karras et al. 2017) and VGGFace2 (Cao et al. 2018). As demonstrated in Table 7 and Figure 12, adopting Impasto in these natural domains can also enhance the quality of protected images without compromising the protection performance of baseline models. This broad applicability highlights Impasto’s versatility and potential as a universal tool for protecting users’ copyright and preventing serious threats of deepfakes.
Appendix E Additional Results
Impasto with other protection methods. In addition to PhotoGuard (Salman et al. 2023) and AdvDM (Liang et al. 2023), we extend the application of Impasto to recent protection techniques, Anti-DreamBooth (Van Le et al. 2023) and Mist (Liang and Wu 2023). Anti-DreamBooth, a variant of AdvDM, incorporates DreamBooth tuning into its optimization process. Mist utilizes a hybrid approach, combining both Encoder (e.g. PhotoGuard) and UNet-based (e.g. AdvDM) protection methods (similar to Eq. 4). We adhere to the hyperparameters specified in their official implementations. Regarding Impasto, we employ the same settings used in our application to AdvDM (Section B), for both Anti-DreamBooth and Mist. As shown in Table 8, Impasto notably enhances the quality of the protected images with comparable protection performances. Figure 13 and 14 also support the efficacy of our method; Impasto significantly diminishes visual artifacts in the protected images while preserving the original methods’ style protection ability. These experiments highlight the adaptability and efficacy of Impasto across various protection frameworks.
Qualitative results. Figure 15 and 16 provide additional visual comparisons, illustrating the impact of applying Impasto in conjunction with PhotoGuard (Salman et al. 2023) and AdvDM (Liang et al. 2023), respectively.
Robustness and generalization. We compare the robustness on countermeasure of Impasto in Figure 17, including JPEG compression, blurring, and noise addition. Across all these countermeasures, Impasto demonstrates a protection capability comparable to that of the baseline. In Figure 18, we present additional results demonstrating the generalization capability of Impasto. When applied to three different personalization methods—DreamBooth, LoRA, and DreamStyler—Impasto maintains performance levels akin to those observed in their respective non-Impasto implementations. This highlights the effectiveness and adaptability of Impasto across diverse personalization techniques.
@@ Appendix E Additional Results
Figure 13: Qualitative comparison of Anti-Dreambooth (Van Le et al. 2023) with and without Impasto. Best viewed in zoom.
@@
Figure 14: Qualitative comparison of Mist (Liang and Wu 2023) with and without Impasto. Best viewed in zoom.
Figure 15: Qualitative comparison of PhotoGuard (Salman et al. 2023) with and without Impasto. Best viewed in zoom.
Figure 16: Qualitative comparison of AdvDM (Liang et al. 2023) with and without Impasto. Best viewed in zoom.
Baseline Refer to caption Refer to caption Refer to caption
+ Impasto Refer to caption Refer to caption Refer to caption
Baseline Refer to caption Refer to caption Refer to caption
+ Impasto Refer to caption Refer to caption Refer to caption
    JPEG Blur Noise
Figure 17: Qualitative comparison on robustness. Methods with Impasto exhibit comparable protection to the baselines.
Baseline Refer to caption Refer to caption Refer to caption Refer to caption
+ Impasto Refer to caption Refer to caption Refer to caption Refer to caption
Baseline Refer to caption Refer to caption Refer to caption Refer to caption
+ Impasto Refer to caption Refer to caption Refer to caption Refer to caption
    Original DreamBooth LoRA DreamStyler
Figure 18: Qualitative comparison on generalization. Impasto does not impede baseline’s generalization abilities across diverse personalization methods; DreamBooth (Ruiz et al. 2023), LoRA (Hu et al. 2021), and DreamStyler (Ahn et al. 2023)

### v4

@@ IV-B Model Analysis
TABLE IX: Quantitative comparison on cartoon and object datasets.
Dataset Method Image Quality Prot. performance
DISTS ($\downarrow$) FID ($\uparrow$)
iCartoonFace PhotoGuard 0.242 (+0.000) 159.4
+ Impasto 0.205 (+0.037) 158.2
AdvDM 0.234 (+0.000) 156.3
+ Impasto 0.211 (+0.023) 157.9
SOP dataset PhotoGuard 0.274 (+0.000) 248.3
+ Impasto 0.258 (+0.016) 250.2
AdvDM 0.215 (+0.000) 281.5
+ Impasto 0.199 (+0.016) 277.2
V Conclusion
We propose Impasto to prevent style imitation with a perceptual-oriented approach. By incorporating perceptual and difficulty-aware protections, along with a perceptual constraint bank, Impasto significantly enhances the quality of protected images within existing protection frameworks. Through a series of benchmarks, we demonstrate that applying Impasto to baseline protection methods improves imperceptibility without compromising protection performance or robustness.
Limitations. All the current protections mostly adopt adversarial perturbations, which can hamper usability due to the extensive time required for the optimization. Even accessible software [53] takes 30-60 mins to protect a $512^{2}$ image on CPU. Addressing this time constraint is a challenge that future research should aim to overcome.
Appendix A Method Details
A-A JND Estimations
  • Luminance adaptation (LA): Perturbations are less visible in regions of very low or high luminance and more noticeable in moderate lighting conditions [20]. Hence, we modulate protection strength based on pixel luminance with a fixed adaptation model.
  • Contrast masking (CM): Perturbations can seamlessly blend into regions with intricate textures, while they leave distinct traces on flat surfaces. To simulate this, we utilize the luminance contrast (or change) of a region to measure the complexity of the pixel [26, 63].
  • Contrast sensitivity function (CSF): Given the band-pass characteristics of the human visual system, we utilize a frequency-based JND model [62]. The human eye is receptive to signals at modulated frequencies while exhibiting insensitivity to high-frequency components. Consequently, perturbations overlaid on high-frequency signals (e.g. edges) are less perceptible.
  • Standard deviation: To assess the spatial structure of an image, we calculate the standard deviation of local image blocks, inspired by SSIM [61]. This measures the image’s structural complexity, which correlates with the sensitivity to subtle perturbations.
  • Entropy: The entropy of an image block is computed to quantify the amount of information or complexity within a local region [64].
Luminance adaptation. The visibility thresholds within the human visual system vary to luminance levels [20]. For instance, our eye is more sensitive under moderate lighting conditions, while discrimination is challenging in a very dark light. Similarly, we observed that perturbations are less noticeable in regions with extremely low or high luminance, prompting an increase in protection strength within these areas. To implement this concept, luminance adaptation is computed following Chou and Li [5] as in the below equation.
[…]
Contrast masking. It is obvious that stimuli become less perceptible against patterned, non-uniform backgrounds. In our context, perturbations can seamlessly blend into regions with intricate textures, while they leave distinct traces on flat surfaces. To simulate this effect, we first compute the luminance contrast [26] by convolving an input image with four directional filters. This process highlights the complexity of a region by contrasting it with its immediate surroundings. Subsequently, contrast masking is established by mapping the luminance contrast onto a logarithmic curve [63].
where luminance contrast $LC(\mathbf{x})$ is obtained via four directional filters as $LC(\mathbf{x})=\frac{1}{16}\times\text{max}_{k=1,\dots,4}|\mathbf{x}*\nabla_{k}|$ following Wu et al. [65].
Contrast sensitivity function. The human visual system exhibits a band-pass response to spatial frequency and CSF represents a function of how our eye is sensitive to the contrast of signals at various spatial frequencies. High spatial frequencies correspond to rapid changes in image details, and contrast sensitivity is optimal at moderate spatial frequencies. Beyond a certain frequency threshold, known as the resolution limit, our eyes are unable to detect changes. Upon this concept, several studies have introduced and modified the CSF [25, 36, 8, 34]; the CSF model $H(f,\theta)$ is given below.
where $\alpha=0.0192,\beta=0.114$ and $f$ is the radial spatial frequency, measured in cycles per degree of visual angle (c/deg). The variable $\theta$, ranging from $\pi,\pi]$, denotes the orientation. Additionally, $f_{\theta}=f/[0.15\times\text{cos}(4\theta)+0.85]$ accounts for the oblique effect [25].
The CSF model is now applied to an image in the frequency domain as $\mathbf{x}_{csf}=\mathcal{F}^{-1}[H(u,v)\times\mathcal{F}(\check{\mathbf{x}})]$, where $\mathcal{F}[\cdot]$ and $\mathcal{F}^{-1}[\cdot]$ represent the DFT and its inverse, respectively. Here, $H(u,v)$ is the DFT-version of $H(f,\theta)$, with $u,v$ being the DFT indices. The transformation from $H(f,\theta)$ to $H(u,v)$ is elaborated in Larson et al. [25]. Prior to applying the CSF model, an input image $\mathbf{x}$ is converted into a perception-adjusted luminance form to reflect the non-linear relationship between digital pixel values and physical luminance [25]. This is achieved by calibrating $\mathbf{x}$ to the settings of an sRGB display and converting it into perceived luminances, which indicates the relative lightness: $\check{\mathbf{x}}=\sqrt[3]{(0.02874\mathbf{x})^{2.2}}$.
Standard deviation. It has served as a crucial metric for quantifying the structural information of an image, both in pixel space [61] and feature space [10]. Given that perturbations tend to be less noticeable in areas exhibiting high levels of change, (as discussed in contrast masking), we focus on these sudden changes, interpreting them as structural components. For this purpose, we compute the block-wise standard deviation in pixel space using a $9\times 9$ local block.
Post-processing All the JND estimations are min-max normalized and then inverted by subtracting from one. When constructing a perceptual map $\mathcal{M}$ with JND estimations, we encountered issues with some JNDs displaying extremely skewed distributions. Additionally, the distributions of JNDs often do not align with each other, posing challenges in their combination. While standardization aligns the distributions, we found that the continuous JND values provide weak signals as masks (e.g. on average, the protection strength drops to 50% compared to the original). To give a more distinct signal depending on the JND values, we discretize the scores by quantizing the JND values. This involves calculating JND quantiles, where the first quantile is set to 1.0. For each subsequent quantile, we multiply by a factor of $\beta$ to decrease the value in a discrete manner. We use $\beta=0.85$, resulting in quantized JND values across four quantiles as [1.0, 0.85, 0.7225, 0.6141]. The value of $\beta$ was selected via grid search on the validation set to achieve the best trade-off between protection performance and imperceptibility. This allows for a more nuanced adjustment of the protection intensity while accommodating the varied distributions of JND estimations.
A-B Perceptual Constraints
Masked LPIPS. We use AlexNet [24] as the backbone network for the LPIPS loss, adhering to the parameter settings in the official LPIPS implementation.
Masked low-pass. In this constraint, we utilize $\text{LP}(\cdot)$, a reconstruction function focusing on the low-frequency component. Inspired by Luo et al. [32], we implement a DWT-based reconstruction module. Given an input image $\mathbf{x}$, DWT decomposes it into one low-frequency component and three high-frequency components, as expressed by the following equation.
$\mathbf{x}_{ll}$ $=\mathbf{LxL}^{T},\;\;\mathbf{x}_{lh}=\mathbf{HxL}^{T},$ (18)
$\mathbf{x}_{hl}$ $=\mathbf{LxH}^{T},\;\;\mathbf{x}_{hh}=\mathbf{HxH}^{T}.$
Appendix B Experimental Settings
Datasets. In our study, we chose the painting and cartoon domains, which are most significantly impacted by style imitation from diffusion models, and used them as benchmark datasets. To create the “painting” dataset, we gather artworks from the WikiArt dataset [58]. This dataset includes selections from 15 artists, 10 artworks per each. The “cartoon” dataset is sourced from the NAVER WEBTOON platform; images are cropped to contain only character depictions. This comprises 15 cartoons with 10 cartoon character images each. Following previous dataset preparation protocols [59], each image was resized to $\times512$. And note that as Impasto does not require a separate training process, the data was used as benchmark set.
TABLE X: Quantitative comparison of protection methods with and without Impasto. Unlike Table I, we match the protected image quality to a similar level and compare the protection performance.
@@ Appendix B Experimental Settings
+ Impasto 0.186 0.346 0.889 4.451 (+0.145) 22.02 (+1.03) 286.9 (+9.3)
AdvDM 0.167 0.730 0.846 3.761 (+0.000) 12.45 (+0.00) 269.0 (+0.0)
+ Impasto 0.164 0.503 0.842 3.947 (+0.231) 11.60 ( - 0.85) 276.6 (+7.6)
+ Impasto 0.241 0.894 0.779 5.480 (+0.443) 12.88 (+2.69) 162.5 (+6.6)
AdvDM 0.241 0.776 0.775 4.802 (+0.000) 10.95 (+0.00) 153.5 (+0.0)
+ Impasto 0.243 0.819 0.744 4.780 ( - 0.022) 11.90 (+0.95) 155.9 (+2.4)
Appendix C Additional Results
@@
Protection performance comparison. Here, unlike Table I, which compares the protected images’ quality under similar protection performance, we compare the protection performance while fixing the image quality. In Table X, Impasto improves protection performance compared to the baseline in most metrics and scenarios. This was also shown in the quality-protection trade-off (Figure 7). These results indicate that attaching Impasto to the protection baselines enables them to be better protection methods.
Qualitative results. Fig. 16 and 17 provide visual comparisons with and without Impasto on Anti-DreamBooth [59] and Mist [30], respectively.
@@
Figure 16: Qualitative comparison of Anti-Dreambooth [59] with and without Impasto. Best viewed in zoom.
@@
Figure 17: Qualitative comparison of Mist [30] with and without Impasto. Best viewed in zoom.